The price of a commodity falls from Rupees 50 to Rupees 30 resulting in an increase in the purchase of the commodity from 200 to 220 units. Calculate the price elasticity of demand.


1
Expert's answer
2021-06-16T04:26:34-0400

The price elasticity of demand is:

"Ed = \frac{220 - 200} {30 - 50} ×\frac{30 + 50} {220 + 200} = -4/21,"

so the demand is inelastic.